# Video ReCap: Recursive Captioning of Hour-Long Videos

### v3

@@ Abstract
Most video captioning models are designed to process short video clips of few seconds and output text describing low-level visual concepts (e.g., objects, scenes, atomic actions). However, most real-world videos last for minutes or hours and have a complex hierarchical structure spanning different temporal granularities. We propose Video ReCap, a recursive video captioning model that can process video inputs of dramatically different lengths (from 1 second to 2 hours) and output video captions at multiple hierarchy levels. The recursive video-language architecture exploits the synergy between different video hierarchies and can process hour-long videos efficiently. We utilize a curriculum learning training scheme to learn the hierarchical structure of videos, starting from clip-level captions describing atomic actions, then focusing on segment-level descriptions, and concluding with generating summaries for hour-long videos. Furthermore, we introduce Ego4D-HCap dataset by augmenting Ego4D with 8,267 manually collected long-range video summaries. Our recursive model can flexibly generate captions at different hierarchy levels while also being useful for other complex video understanding tasks, such as VideoQA on EgoSchema. Data, code, and models are available [1].
@@ 2 Related Works
Video Captioning Methods. Early works in video captioning used template-based approaches [27, 25, 43, 49, 61]. Subsequently, these methods were replaced by deep learning methods built using CNN-RNN encoder-decoder architectures [64, 17, 55, 37, 38, 8, 47, 56]. The recent introduction of Transformer [53, 18] led to a plethora of transformer-based video captioning methods [39, 22, 38, 8, 47, 56, 28, 63, 51, 46]. Though these approaches have shown great success in short clip captioning, most are limited to short video clip sequences of a few seconds and, thus, cannot generate captions spanning multiple temporal hierarchies for hour-long videos.
@@ 4 Ego4D-HCap Dataset
While Ego4D comes with time-stamped atomic captions and video-segment descriptions spanning up to 5 minutes, it lacks video-level summaries for longer video durations. To address this issue, we annotate a subset of 8,267 Ego4D videos with long-range video summaries, each spanning up to two hours. This enhancement provides a three-level hierarchy of captions, making it a perfect resource for validating the effectiveness of our model on the hierarchical video captioning task. In Table 1, we provide a detailed summary of our introduced Ego4D-HCap subset.
@@ 5 Experimental Setup
Video ReCap TSF-B GPT2 339M 98.35 48.77 28.28
Video ReCap-U TSF-B GPT2 113M 92.67 47.90 28.08
@@ C R M C R M
Video ReCap TSF-B GPT2 339M 41.74 39.04 18.21 28.06 32.27 14.26
Video ReCap TSF-B GPT2 339M 46.88 39.73 18.55 29.34 32.64 14.45
Video ReCap-U TSF-B GPT2 113M 45.60 39.33 18.17 31.06 33.32 14.16
@@ 5.1 Hierarchical Video Captioning Baselines
      BLIP2 [29]: A zero-shot baseline for short-term clip captioning that utilizes a state-of-the-art image captioning model.
@@
      BLIP2 + GPT3.5 [29, 12]: A zero-shot text-based baseline for video segment descriptions and long-range video summaries. Given BLIP2-generated captions, it uses GPT3.5 to generate video segment descriptions and long-range video summaries.
@@
      LaViLa + GPT3.5 [67, 12]: Similar to the above, a zero-shot baseline for video segment and summary generation using LaViLa captions fed into GPT3.5.
@@
      LaViLa + GPT2 [67, 41]: A fully-finetuned text-based baseline that takes LaViLa-generated clip captions and finetunes a text-only GPT2 model for segment description and video summary generation while keeping the underlying LaViLa model frozen.
@@
      LaViLa + FLAN-T5 [67, 15]: Similar to the above, a fully-finetuned text-based baseline that uses FLAN-T5 rather than GPT2 for segment description and video summary generation.
@@
      LaViLa [67]: A video-based baseline, finetuned end-to-end to generate short-term captions, medium-length segment descriptions, and long-range video summaries directly using video inputs. Note that this baseline uses the same video encoder, text decoder, and other experimental settings as our model.
@@ 6.2 Long-Range VideoQA on EgoSchema
Table 3: Long-Range VideoQA on EgoSchema [35] Our approach achieves state-of-the-art results, outperforming the previous best method, InternVideo, by a substantial margin of 18.13%. Furthermore, leveraging the hierarchical captions produced by our model leads to 4.2% and 5.96% boost in performance compared to using our model’s short-clip captions or captions generated by LaViLa [67]. This demonstrates the efficacy of hierarchical video captions for long-range video question answering.
@@ 6.3 Ablation Studies
Recursive Input Segment Description Video Summary
@@ C R M C R M
40.17 38.65 17.59 25.64 29.61 13.57
41.74 39.04 18.21 28.06 32.27 14.26
Table 4: Inportance of Recursive Inputs. A non-recursive variant of our model performs worse in segment description and video summary generation (-1.57% and -2.42% in CIDEr).
Importance of Recursive Architecture. In this section, we analyze the significance of the recursive architecture of our Video ReCap model. For this variant, we discard the recursive inputs (i.e., captions generated in the previous hierarchy level) and compare the performance with our recursive model. Table 4 shows the result. We observe that the performance of a non-recursive variant drops by 1.57% CIDEr for video segment descriptions. Moreover, the recursive model structure is even more important for long-range video summary generation, where the performance of a non-recursive variant drops by 2.42% CIDEr without it. These experiments reveal that the recursive design of Video ReCap leads to better performance on the hierarchical video captioning task, particularly on long-range video inputs.
Training Scheme Segment Descrption Video Summary
@@
Significance of Hierarchical Curriculum Learning Next, we investigate the significance of our hierarchical curriculum learning scheme. Table 5 shows the importance of such a curriculum learning scheme. We observe that if we directly train our model on the segment description from GPT2 pretrained initialization, performance drops by a significant margin of 4.93% CIDEr. Moreover, the performance drop is even more catastrophic (-19.44%) for video summary generation without curriculum learning. Finally, we show that it is useful to progressively incorporate higher-level captions, starting from short-term captions, then transitioning to medium-length segment descriptions, and lastly, finishing with long-range video summaries. The variant that progresses from short-term caption to long-range video summary learning directly exhibits a 3.22% drop in CIDEr performance.
@@
Importance of LLM-Based Supervision Next, we study the importance of LLM-based supervision for medium-length segment descriptions and long-range video summaries. In Table 5(a), we show the performance of different LLM Teachers (e.g., GPT2 [41], and FLAN-T5 [15]) that we use to generate the pseudo ground truth data. We observe that FLAN-T5-Large achieves the best performance in all metrics. Hence, we use FLAN-T5-Large as our Teacher to generate pseudo-ground truth data for segment descriptions and long-range video summaries. Specifically, we produce 100K pseudo-annotations for segment descriptions and 15K for video summaries. We combine these pseudo-annotations with the manually annotated data and train our model. Table 5(b) shows that utilizing supervision from LLMs provides a substantial performance boost in both segment description (+5.14% CIDEr gain) and video summary (+1.28% CIDEr improvement) generation performance.
Input Segment Description Video Summary
C R M C R M
Video 40.17 38.65 17.59 25.64 29.61 13.57
Text 40.10 38.02 17.41 23.23 29.17 13.31
Video + Text 41.74 39.04 18.21 28.06 32.27 14.26
Table 7: Video-Language Input Ablation. Using both video and text features leads to better performance for both segment description and video summary generation.
Ablation of Input Modalities. As described in Section 3.2, our model utilizes both video features and text inputs (generated in the previous hierarchy) for the segment descriptions and video summaries. Note that we do not use any text inputs for clip captions as they define the base case of our recursive video model. Since we need to sparsely sample video features to fit long-range videos into GPU memory, we hypothesize that using text as an intermediate representation should complement the sparse video features. Table 7 proves our hypothesis and shows that using both video and text features as inputs yields the best performance for our model. Specifically, for segment description generation, combining video and text inputs produces a +1.57% boost relative to video-only and +1.64% boost relative to text-only baselines in CIDEr. Moreover, for long-range video summary generation, video + text inputs provide +2.42% and +4.83% gains compared to video-only and text-only variants.
@@ 7 Conclusions and Future Work
Acknowledgements. We thank Feng Cheng, Yan-Bo Lin, Ce Zhang, Yue Yang, and Soumitri Chattopadhyay for helpful discussions. This work was supported by the Sony Faculty Innovation award, Laboratory for Analytic Sciences via NC State University, ONR Award N00014-23-1-2356.
@@
Our supplementary materials contain Section S1: Additional Implementation Details, Section S2: Ego4D-HCap Data Collection Process, Section S3: Additional Quantitative Results, and Section S4: Qualitative Results.
@@
S2.4 Example Video Summaries.
@@
S3 Additional Quantitative Results
@@ C R M C R M C R M
S4 Qualitative Results on EgoSchema

### v4

@@ Abstract
Most video captioning models are designed to process short video clips of few seconds and output text describing low-level visual concepts (e.g., objects, scenes, atomic actions). However, most real-world videos last for minutes or hours and have a complex hierarchical structure spanning different temporal granularities. We propose Video ReCap, a recursive video captioning model that can process video inputs of dramatically different lengths (from 1 second to 2 hours) and output video captions at multiple hierarchy levels. The recursive video-language architecture exploits the synergy between different video hierarchies and can process hour-long videos efficiently. We utilize a curriculum learning training scheme to learn the hierarchical structure of videos, starting from clip-level captions describing atomic actions, then focusing on segment-level descriptions, and concluding with generating summaries for hour-long videos. Furthermore, we introduce Ego4D-HCap dataset by augmenting Ego4D with 8,267 manually collected long-range video summaries. Our recursive model can flexibly generate captions at different hierarchy levels while also being useful for other complex video understanding tasks, such as VideoQA on EgoSchema. Data, code, and models are publicly available at https://sites.google.com/view/vidrecap.
@@ 2 Related Works
Video Captioning Methods. Early works in video captioning used template-based approaches [26, 24, 42, 48, 60]. Subsequently, these methods were replaced by deep learning methods built using CNN-RNN encoder-decoder architectures [63, 16, 54, 36, 37, 7, 46, 55]. The recent introduction of Transformer [52, 17] led to a plethora of transformer-based video captioning methods [38, 21, 37, 7, 46, 55, 27, 62, 50, 45]. Though these approaches have shown great success in short clip captioning, most are limited to short videos of a few seconds and cannot generate captions spanning multiple temporal hierarchies for hour-long videos.
@@ 4 Ego4D-HCap Dataset
While Ego4D comes with time-stamped atomic captions and video-segment descriptions spanning up to 5 minutes, it lacks video-level summaries for longer video durations. To address this issue, we annotate a subset of 8,267 Ego4D videos with long-range video summaries, each spanning up to two hours. This enhancement provides a three-level hierarchy of captions, making it a perfect resource for validating the effectiveness of our model on the hierarchical video captioning task. In Table 1, we provide a detailed summary of our introduced Ego4D-HCap subset. Please refer to our supplementary materials for a more detailed analysis of the Ego4D-HCap dataset.
@@ 5 Experimental Setup
HierVidCap TSF-B GPT2 339M 98.35 48.77 28.28
HierVidCap-U TSF-B GPT2 113M 92.67 47.90 28.08
@@ C R M C R M
HierVidCap TSF-B GPT2 339M 41.74 39.04 18.21 28.06 32.27 14.26
HierVidCap TSF-B GPT2 339M 46.88 39.73 18.55 29.34 32.64 14.45
HierVidCap-U TSF-B GPT2 113M 45.60 39.33 18.17 31.06 33.32 14.16
@@ 5.1 Hierarchical Video Captioning Baselines
      BLIP2 [28]. A zero-shot baseline for short-term clip captioning that utilizes a state-of-the-art image captioning model.
@@
      BLIP2 + GPT3.5 [28, 11]. A zero-shot text-based baseline for video segment descriptions and long-range video summaries. Given BLIP2-generated captions, it uses GPT3.5 to generate video segment descriptions and long-range video summaries.
@@
      LaViLa + GPT3.5 [66, 11]. Similar to the above, a zero-shot baseline for video segment and summary generation using LaViLa captions fed into GPT3.5.
@@
      LaViLa + GPT2 [66, 40]. A fully-finetuned text-based baseline that takes LaViLa-generated clip captions and finetunes a text-only GPT2 model for segment description and video summary generation while keeping the underlying LaViLa model frozen.
@@
      LaViLa + FLAN-T5 [66, 14]. Similar to the above, a fully-finetuned text-based baseline that uses FLAN-T5 rather than GPT2 for segment description and video summary generation.
@@
      LaViLa [66]. A video-based baseline, finetuned end-to-end to generate short-term captions, medium-length segment descriptions, and long-range video summaries directly using video inputs. Note that this baseline uses the same video encoder, text decoder, and other experimental settings as our model.
@@ 6.2 Long-Range VideoQA on EgoSchema
Table 3: Long-Range VideoQA on EgoSchema [34] Our approach achieves state-of-the-art results, outperforming the previous best method, InternVideo, by a substantial margin of 18.13%. Furthermore, leveraging the hierarchical captions produced by our model leads to 4.2% and 5.96% boost in performance compared to short-clip captions generated by Video ReCap or LaViLa [66].
@@ 6.3 Ablation Studies
Ablation of Input Modalities. Our model utilizes both video features and recursive text inputs (generated in the previous hierarchy) for the segment descriptions and video summaries. Note that we do not use any text inputs for clip captions as they define the base case of our recursive video model. Since we need to sparsely sample video features to fit long-range videos into GPU memory, we hypothesize that using text as an intermediate representation should complement the sparse video features. In Table 4, we compare our model with a non-recursive baseline (row 2), which only uses sparse video features and a recursive baseline (row 3), which only uses recursive text features. We observe that combining video and text inputs produces a +1.57% boost relative to video-only and a +1.64% boost compared to text-only baselines in CIDEr for segment description generation. Moreover, combining both inputs is more important for long-range video summary generation, where video+text inputs provide +2.42% and +4.83% gains compared to video-only and text-only variants. These experiments reveal that the recursive design of Video ReCap that utilizes both video and text input modalities is crucial for the hierarchical video captioning task.
Input Segment Description Video Summary
@@ C R M C R M
Video 40.17 38.65 17.59 25.64 29.61 13.57
Text 40.10 38.02 17.41 23.23 29.17 13.31
Video + Text 41.74 39.04 18.21 28.06 32.27 14.26
Table 4: Video-Language Input Ablation. Using both sparse video features and recursive text inputs leads to better performance for both segment description and video summary generation.
Training Scheme Segment Description Video Summary
@@
Significance of Hierarchical Curriculum Learning. Next, we investigate the significance of our hierarchical curriculum learning scheme. Table 5 shows the importance of such a curriculum learning scheme. We observe that if we directly train our model on the segment description from GPT2 pretrained initialization, performance drops by a significant margin of 4.93% CIDEr. Moreover, the performance drop is even more catastrophic (-19.44%) for video summary generation without curriculum learning. Finally, we show that it is useful to progressively incorporate higher-level captions, starting from short-term captions, then transitioning to medium-length segment descriptions, and lastly, finishing with long-range video summaries. The variant that progresses from short-term caption to long-range video summary learning directly exhibits a 3.22% drop in CIDEr performance.
@@
Importance of LLM-Based Supervision. Finally, we study the importance of LLM-based supervision for medium-length segment descriptions and long-range video summaries. In Table 6(a), we show the performance of different LLM Teachers (e.g., GPT2 [40], and FLAN-T5 [14]) that we use to generate the pseudo ground truth data. We observe that FLAN-T5-Large achieves the best performance in all metrics. Hence, we use FLAN-T5-Large as our Teacher to generate pseudo-ground truth data for segment descriptions and long-range video summaries. Specifically, we produce 100K pseudo-annotations for segment descriptions and 15K for video summaries. We combine these pseudo-annotations with the manually annotated data and train our model. Table 6(b) shows that utilizing supervision from LLMs provides a substantial performance boost in both segment description (+5.14% CIDEr gain) and video summary (+1.28% CIDEr improvement) generation performance.
@@ 7 Conclusions and Future Work
Acknowledgements. We thank Feng Cheng, Yan-Bo Lin, Ce Zhang, Yue Yang, and Soumitri Chattopadhyay for their helpful discussions. Authors from UNC Chapel Hill were supported by the NIH Award R01HD11107402, Sony Faculty Innovation award, Laboratory for Analytic Sciences via NC State University, and ONR Award N00014-23-1-2356.
@@
Our supplementary materials contain Section S1: Additional Implementation Details, Section S2: Ego4D-HCap Data Collection Process, Section S3: Ego4D-HCap Dataset Analysis, Section S4: Additional Quantitative Results, and Section S5: Qualitative Results on EgoSchema.
@@
S3 Ego4D-HCap Dataset Analysis
S3.1 Example Video Summaries.
@@
S4 Additional Quantitative Results
@@ C R M C R M C R M
S5 Qualitative Results on EgoSchema